# Conformal Validity Guarantees Exist for Any Data Distribution

### v2

@@ Abstract
As machine learning (ML) gains widespread adoption, practitioners are increasingly seeking means to quantify and control the risk these systems incur. This challenge is especially salient when ML systems have autonomy to collect their own data, such as in black-box optimization and active learning, where their actions induce sequential feedback-loop shifts in the data distribution. Conformal prediction has emerged as a promising approach to uncertainty and risk quantification, but prior variants’ validity guarantees have assumed some form of “quasi-exchangeability” on the data distribution, thereby excluding many types of sequential shifts. In this paper we prove that conformal prediction can theoretically be extended to any joint data distribution, not just exchangeable or quasi-exchangeable ones, although it is exceedingly impractical to compute in the most general case. For practical applications, we outline a procedure for deriving specific conformal algorithms for any data distribution, and we use this procedure to derive tractable algorithms for a series of ML-agent-induced covariate shifts. We evaluate the proposed algorithms empirically on synthetic black-box optimization and active learning tasks.
@@ 2.1 Standard Conformal Prediction
For an accessible introduction to CP see Angelopoulos et al. (2022) or for the definitive reference see Vovk et al. (2005). Computing conformal predictive sets requires access to a real-valued score function $\mathcal{S}$ that takes as its inputs a point $(x,y)$ and a “bag” [^3] That is, a multiset. CP methods assume the score-fitting algorithm is symmetric with respect to permutations of its training inputs $\bar{Z}$; we maintain this assumption unless stated otherwise. of other examples $\bar{Z}$, where $\mathcal{S}\big{(}(x,y),\bar{Z}\big{)}$ can be thought of as quantifying how “nonconforming” or “strange” the point $(x,y)$ is relative to $\bar{Z}$. For example, the absolute value residual $|\widehat{\mu}_{\bar{Z}}(x)-y|$ is a common score function, where $\widehat{\mu}_{\bar{Z}}$ is an ML predictor trained on the examples $\bar{Z}$. For shorthand, we will sometimes write the fitted score function as $\widehat{\mathcal{S}}(x,y)=\mathcal{S}\big{(}(x,y),\bar{Z}\big{)}$ and outputted score values as $V_{i}=\widehat{\mathcal{S}}(Z_{i})$. A standard CP set $\widehat{\mathcal{C}}_{n}(x)$ is then constructed by taking a (conservative) quantile on the empirical distribution of score values for each candidate $y\in\mathcal{Y}$
$\widehat{\mathcal{C}}_{n}(x)=\Big{\{}y\in\mathcal{Y}:\widehat{\mathcal{S}}(x,y)\leq\mathrm{Q}_{1-\alpha}\big{(}V_{1:n}\cup\{\infty\}\big{)}\Big{\}},$ (2)
where $Q_{1-\alpha}$ is an empirical quantile evaluated at level $1-\alpha$.
Standard CP methods assume that all the data $Z_{1},\dots,Z_{n+1}$ are exchangeable—that is, that their joint distribution is invariant to permutations, a special case being independent and identically distributed (IID). Under exchangeability, $\widehat{\mathcal{C}}_{n}(x)$ in Eq. (2) achieves the valid coverage guarantee in Eq. (1) with finite samples. This standard CP guarantee is distribution-free in the sense of making no assumptions about the specific parametric form of the data distribution—this is not to be confused with robustness to distribution shift, which violates exchangeability for example when the distribution of $Z_{n+1}$ differs from that of $Z_{1:n}$. Exchangeability is thus still a strong assumption in its own way. For instance, it excludes many environment-dependent or time-dependent distribution shifts which often occur in practice.
@@ 2.2 Weighted Conformal Prediction for Covariate Shift
Note that in standard covariate shift the test distribution $P_{Z}^{(1)}$ is independent from the training data; it is fixed a priori and cannot change depending on different draws of $Z_{1:n}$. Tibshirani et al. (2019)’s weighted CP algorithm for standard covariate shift generalizes Eq. (2) with likelihood ratio function weights $w(x)=dP_{X}^{(1)}(x)/dP_{X}^{(0)}(x)$.
Fannjiang et al. (2022) followed the general approach of Tibshirani et al. (2019) to derive a specific weighted CP algorithm with a valid coverage guarantee (Eq. (1)) in a setting they call (one-step) feedback covariate shift (FCS), where $P_{X;Z_{1:n}}^{(1)}$ is explicitly dependent on the realized draw $Z_{1:n}$. While this work was an important extension of CP to dependent data, they did not consider a more general “multistep” FCS case that could describe how data shifts may accumulate over time, for instance in active learning. Their focus on the one-step FCS case could be because they posit as the basis of their theoretical analysis an assumption they call pseudo-exchangeability, which explicitly characterizes one-step FCS but does not formally describe its multistep analog (see Appendix C.3), which we introduce next.
@@ 2.3 Multistep Feedback Covariate Shift
$(X_{n+t},Y_{n+t})$ $\hskip 1.42262pt\sim P_{X;Z_{1:(n+t-1)}}^{(t)}\times P_{Y\mid X},\;t=1,\dots,T.$
Note that $n$ is the initial sample size of an IID initialization, and $t$ is the number of MFCS steps past this initialization. Setting $T=1$ reduces Eq. (3) to one-step FCS (Fannjiang et al., 2022) and moreover to standard covariate shift if independence is further assumed; we also note that MFCS is similar to “sequential ignorability” in causal inference.
MFCS is commonly induced by ML agents whenever the ML system is able to actively select the next datapoint $X_{n+t}$ (e.g., based on a cost or utility criterion), informed by its prior experience $Z_{1:(n+t-1)}$, so long as the predictive goal or “concept” of interest (i.e., $Y\mid X$) remains the same. For instance, MFCS describes active learning, where the goal is to efficiently construct a training corpus to improve model performance, as well as black-box optimization, where the ML model is an instrument to optimize $X$ with respect to some unknown cost (or utility) function that is only indirectly accessible by each observation of $Y$.
@@ 3 Related Work
Conformal prediction under distribution shift. Prior work can largely be categorized into variants of weighted CP (Tibshirani et al., 2019; Podkopaev & Ramdas, 2021; Xu & Xie, 2021; Fannjiang et al., 2022; Prinster et al., 2022, 2023; Barber et al., 2023; Farinhas et al., 2023; Nair & Janson, 2023; Yang et al., 2024) and adaptive CP (Gibbs & Candès, 2021, 2022; Zaffran et al., 2022; Angelopoulos et al., 2023; Feldman et al., 2023; Bhatnagar et al., 2023). The former exploits knowledge or estimates of the data shift to proactively adjust inference and thereby achieve guarantees that hold at inference time; meanwhile, the latter aims to maintain a target risk threshold by retroactively responding to threshold violations and thus obtain guarantees most meaningful “in the long run” rather than at inference time.
@@
Conformal inference for decision-making. Ideas from CP are increasingly being applied to decision-making to inform utility estimates (Vovk & Bendtsen, 2018; Stanton et al., 2023; Salinas et al., 2023; Lekeufack et al., 2023) or satisfy constraints (Zhang et al., 2023; Dixit et al., 2023; Laufer-Goldshtein et al., 2023; Jin & Candès, 2023). In particular, Stanton et al. (2023) drew motivation from Fannjiang et al. (2022) to combine CP with Bayesian optimization but noted that a gap in theory prevented a formal coverage guarantee after the first round, a gap we address in this work.
@@ 4.1 The Role of (Weighted) Exchangeability and Related Assumptions in Conformal Prediction
Informal Statement of the Key Insight One can think of CP as an inverted permutation test (with a conservative adjustment). Informally, the key insight is that—from this perspective of an inverted permutation test—the only role of exchangeability, weighted exchangeability (Tibshirani et al., 2019), or related distributional assumptions such as pseudo-exchangeability (Fannjiang et al., 2022) in CP is to enable tractability by telling us how much to “weigh” or “count” a given ordering of datapoints; in particular, by facilitating computation of the relative likelihood of each sequence of observations, relative to some (permutation-invariant) core function, given the unordered “bag” of observed values (details in Appendix C). So, extensions of CP should be possible so long as one can quantify the probability of any ordering of observed datapoints (or score values).
Formal Sketch of the Key Insight We first sketch analysis from Tibshirani et al. (2019) (which implicitly constructs CP as an inverted permutation test) followed by the key insight for our current paper’s main result. Their analysis begins by conditioning on the set [^4] For simplicity, we assume all the values are distinct to work with sets rather than multisets; in general, the argument holds with more complex notation using uniform randomness to break ties. of data values (or scores)—that is, on the event $\{Z_{1},...,Z_{n+1}\}=\{z_{1},...,z_{n+1}\}$ denoted by $E_{z}$. To be clear, $E_{z}$ means the unordered “bag” of values is known but not whether $Z_{i}=z_{i}$, and so on; we know that the random variable $Z_{i}$ takes on an observed value in $\{z_{1},...,z_{n+1}\}$, but we do not yet know which value (similarly for the scores $V_{i}$). Given $E_{z}$, the goal is to examine the probability that the test point’s score $V_{n+1}$ was actually $v_{i}$, for each $i\in\{1,...,n+1\}$, which they write as
@@
where $f$ denotes the joint probability density function (PDF). Standard CP proceeds from this statement by substituting $f(z_{\sigma(1)},...,z_{\sigma(n+1)})=f(z_{1},...,z_{n+1})$ via exchangeability, reducing the problem to counting permutations; weighted CP proceeds by factorizing $f$ using weighted exchangeability, with the product of weight functions representing the “weight” given to each permutation $\sigma$; and, CP with pseudo-exchangeable data proceeds similarly to handle a specific instance of potential dependencies (details in Appendix C).
Generally, the key insight in our paper is that simplifying Eq. (4) is only a practical rather than a theoretically necessary step. For intuition on this point, note that Eq. (4) makes no assumptions on $f$; it follows from the definition of conditional probability and law of total probability (LOTP):
$\mathbb{P}\{Z_{n+1}=z_{i}\ |\ E_{z}\}$ =$=\frac{\mathbb{P}\{Z_{n+1}=z_{i},E_{z}\}}{\mathbb{P}\{E_{z}\}}$
@@
That is, the last step uses LOTP in both the numerator and denominator—for instance, the denominator’s sum is over every way the event $E_{z}$ could occur, which amounts to all the possible permutations $\sigma$ of the values. The crux of extending CP to any, possibly non-exchangeable $f$ can thus reduce to computing $f(z_{\sigma(1)},...,z_{\sigma(n+1)})$ directly for all orderings $\sigma$, or factoring $f$ into “dyanmic” terms that are tractable to compute (e.g., “weight functions”) and remaining permutation-invariant terms that cancel in the ratio.
@@ 4.2 A General View of Conformal Prediction
Limitations and Benefits of the General View of CP There are two main obstacles to computing Eq. (6) in practice: a computational complexity challenge and an epistemic challenge regarding knowledge about $f$. That is, in the general case for an arbitrary or unfactorized $f$, Eq. (4) requires $\mathcal{O}((n+1)!)$ evaluations of $f$, which quickly becomes intractable. Secondly, we often do not know $f$; indeed, the primary motivation of distribution-free UQ is that very deficiency. Exchangeability, weighted exchangeability (Tibshirani et al., 2019), and related conditions such as pseudo-exchangeability (Fannjiang et al., 2022) introduce distributional assumptions which alleviate these challenges.
@@
One immediate benefit of the general presentation of conformal prediction in Theorem 4.1 is a shift in perspective: because the theorem holds for any joint distribution $f$, for any particular $f$ the question is not whether a CP set with valid coverage exists, it is instead whether we can compute the prediction set in practice. Consequently—from this general perspective—“quasi-exchangeability” definitions are no longer necessary conditions for CP theory; rather, the only role of these definitions is to enable practical tractability by simplifying and eliminating terms from the computation (of Eq. (4))—for example, at one extreme, exchangeability “assumes away” $f$ from the computation entirely. (Appendix C provides further formal details on this point, including how previous CP validity guarantees can be viewed as corollaries to Theorem 4.1.) Specific weighted CP algorithms with corresponding coverage validity guarantees can thus be derived for any (possibly non-exchangeable) distribution. In section 4.3, we outline how to perform such derivations in general and provide the MFCS case as a worked example.
@@
Further Discussion of Tibshirani et al. (2019) This general view of CP is implicit in a close reading of Tibshirani et al. (2019), [^6] We note that the main insight in our paper is also invoked in the following lecture notes (Tibshirani, 2023): https://www.stat.berkeley.edu/~ryantibs/statlearn-s23/lectures/conformal_ds.pdf, which a sharp reader could use to infer Theorem 4.1. However, based on how Tibshirani et al. (2019) has been cited in the CP literature (i.e., primarily as extending CP to standard covariate shift or to weighted exchangeability), we believe this important insight has not been fully appreciated by the community and we hope that our contribution can help correct this oversight. but in that paper the authors did not explicitly state or prove the general result we present in Theorem 4.1. Instead, Tibshirani et al. (2019) premised their main result on their definition of weighted exchangeability, although this condition is not strictly necessary for CP in the broad sense. While we maintain that weighted exchangeability is still a useful definition (see Appendix C.2), subsequent literature has seemingly been led to believe that new definitions of “quasi-exchangeability” are necessary for advancing and delineating the boundaries of CP theory—for example, Fannjiang et al. (2022)’s introduction of pseudo-exchangeability to study the single-step FCS setting. We hope our contribution will lead to wider recognition of the generality of Tibshirani et al. (2019)’s analysis and foster a broadened scope of possibility in the study of CP: in particular, we encourage a shift away from viewing “quasi-exchangeability” conditions as theoretical boundaries to be overcome, and instead towards primarily using such conditions to clearly specify the assumptions granted by a given setting of interest or adopted for practical tractability.
@@
4.3 Deriving Weighted CP Validity Guarantees and Algorithms for MFCS, or Any Data Distribution

### v3

@@ Abstract
As artificial intelligence (AI) / machine learning (ML) gain widespread adoption, practitioners are increasingly seeking means to quantify and control the risk these systems incur. This challenge is especially salient when such systems have autonomy to collect their own data, such as in black-box optimization and active learning, where their actions induce sequential feedback-loop shifts in the data distribution. Conformal prediction is a promising approach to uncertainty and risk quantification, but prior variants’ validity guarantees have assumed some form of “quasi-exchangeability” on the data distribution, thereby excluding many types of sequential shifts. In this paper we prove that conformal prediction can theoretically be extended to any joint data distribution, not just exchangeable or quasi-exchangeable ones. Although the most general case is exceedingly impractical to compute, for concrete practical applications we outline a procedure for deriving specific conformal algorithms for any data distribution, and we use this procedure to derive tractable algorithms for a series of AI/ML-agent-induced covariate shifts. We evaluate the proposed algorithms empirically on synthetic black-box optimization and active learning tasks.
@@ 2.1 Standard Conformal Prediction
For an accessible introduction to CP see Angelopoulos et al. (2022) or for the definitive reference see Vovk et al. (2005). Computing conformal predictive sets requires access to a real-valued score function $\mathcal{S}$ that takes as its inputs a point $(x,y)$ and a “bag” [^3] That is, a multiset. CP methods assume the score-fitting algorithm is symmetric with respect to permutations of its training inputs $\bar{Z}$; we maintain this assumption unless stated otherwise. of other examples $\bar{Z}$, where $\mathcal{S}\big{(}(x,y),\bar{Z}\big{)}$ can be thought of as quantifying how “nonconforming” or “strange” the point $(x,y)$ is relative to $\bar{Z}$. For example, the absolute value residual $|\widehat{\mu}_{\bar{Z}}(x)-y|$ is a common score function, where $\widehat{\mu}_{\bar{Z}}$ is an ML predictor trained on the examples $\bar{Z}$. For shorthand, we will sometimes write the fitted score function as $\widehat{\mathcal{S}}(x,y)=\mathcal{S}\big{(}(x,y),\bar{Z}\big{)}$ and outputted score values as $V_{i}=\widehat{\mathcal{S}}(Z_{i})$. With $Q_{1-\alpha}$ denoting an empirical quantile evaluated at level $1-\alpha$, a standard CP set $\widehat{\mathcal{C}}_{n}(x)$ is then constructed by taking a (conservative) quantile on the empirical distribution of score values for each $y\in\mathcal{Y}$:
$\widehat{\mathcal{C}}_{n}(x)=\Big{\{}y\in\mathcal{Y}:\widehat{\mathcal{S}}(x,y)\leq\mathrm{Q}_{1-\alpha}\big{(}V_{1:n}\cup\{\infty\}\big{)}\Big{\}}.$ (2)
Standard CP methods assume that all the data $Z_{1},\dots,Z_{n+1}$ are exchangeable—that is, that their joint distribution is invariant to permutations, a special case being independent and identically distributed (IID). Under exchangeability, $\widehat{\mathcal{C}}_{n}(x)$ in Eq. (2) achieves the valid coverage guarantee in Eq. (1) with finite data samples. This standard CP guarantee is distribution-free in the sense of making no assumptions about the specific parametric form of the data distribution—this is not to be confused with robustness to distribution shift, which violates exchangeability for example when the distribution of $Z_{n+1}$ differs from that of $Z_{1:n}$. Exchangeability is thus still a strong assumption in its own way. For instance, it excludes many environment-dependent or time-dependent distribution shifts which often occur in practice.
@@ 2.2 Weighted Conformal Prediction for Covariate Shift
Note that in standard covariate shift the test distribution $P_{Z}^{(1)}$ is independent from the training data; it is fixed a priori and cannot change depending on different draws of $Z_{1:n}$. Tibshirani et al. (2019)’s weighted CP algorithm for standard covariate shift generalizes Eq. (2) with likelihood (or density) ratio function weights $w(x)=p_{X}^{(1)}(x)/p_{X}^{(0)}(x)$.
Fannjiang et al. (2022) followed the overall approach of Tibshirani et al. (2019) to derive a specific weighted CP algorithm with a valid coverage guarantee (Eq. (1)) in a setting they called (one-step) feedback covariate shift (FCS), where $P_{X;Z_{1:n}}^{(1)}$ is explicitly dependent on the realized draw $Z_{1:n}$. While this work was an important extension of CP to dependent data, they did not consider a more general “multistep” FCS case that could describe how data shifts may accumulate over time, for instance in active learning. Their focus on the one-step FCS case could be because they posit as the basis of their theoretical analysis an assumption they call pseudo-exchangeability, which explicitly characterizes one-step FCS but does not formally describe its multistep analog (see Appendix D.3), which we introduce next.
@@ 2.3 Multistep Feedback Covariate Shift
$(X_{n+t},Y_{n+t})$ $\hskip 1.42262pt\sim P_{X;Z_{1:(n+t-1)}}^{(t)}\times P_{Y\mid X},\;t=1,\dots,T,$
where $n$ is the size of an IID initialization, and $t$ is the number of subsequent MFCS steps. Setting $T=1$ reduces Eq. (3) to one-step FCS (Fannjiang et al., 2022) and moreover to standard covariate shift if independence is further assumed. MFCS is commonly induced by ML agents able to actively select each datapoint $X_{n+t}$, informed by prior training observations, so long as the predictive goal or “concept” $Y|X$ remains the same. For instance, MFCS describes active learning, where the goal is to efficiently construct a training corpus to improve model performance, and black-box optimization, where the ML model is an instrument to optimize $X$ with respect to some unknown cost (or utility) function only indirectly accessible by each observation of $Y$.
@@ 3 Related Work
Conformal Prediction Under Distribution Shift Prior work can largely be categorized into variants of weighted CP (Tibshirani et al., 2019; Podkopaev & Ramdas, 2021; Xu & Xie, 2021; Fannjiang et al., 2022; Prinster et al., 2022, 2023; Barber et al., 2023; Farinhas et al., 2023; Nair & Janson, 2023; Yang et al., 2024) and adaptive CP (Gibbs & Candès, 2021, 2022; Zaffran et al., 2022; Angelopoulos et al., 2023; Feldman et al., 2023; Bhatnagar et al., 2023). The former exploits knowledge or estimates of the data shift to proactively adjust inference and thereby achieve guarantees that hold at inference time; meanwhile, the latter aims to maintain a target risk threshold by retroactively responding to threshold violations and thus obtain guarantees most meaningful “in the long run” rather than at inference time.
@@
Conformal Prediction for Decision Making Ideas from CP are increasingly being applied to decision-making to inform utility estimates (Vovk & Bendtsen, 2018; Stanton et al., 2023; Salinas et al., 2023) or satisfy constraints (Lekeufack et al., 2023; Zhang et al., 2023; Dixit et al., 2023; Laufer-Goldshtein et al., 2023; Jin & Candès, 2023). In particular, Stanton et al. (2023) drew motivation from Fannjiang et al. (2022) to combine CP with Bayesian optimization but noted that a gap in theory prevented a formal coverage guarantee after the first round, a gap we address in this work.
@@ 4.1 The Role of (Weighted) Exchangeability and Related Assumptions in Conformal Prediction
Informal Statement of the Key Insight One can think of CP as an inverted permutation test (with a conservative adjustment). Informally, the key insight is that—from this perspective of an inverted permutation test—the only role of exchangeability, weighted exchangeability (Tibshirani et al., 2019), or related distributional assumptions such as pseudo-exchangeability (Fannjiang et al., 2022) in CP is to enable tractability by telling us how much to “weigh” or “count” a given ordering of datapoints; in particular, by facilitating computation of the relative likelihood of each sequence of observations, relative to some (permutation-invariant) core function, given the unordered “bag” of observed values (details in Appendix D). So, extensions of CP should be possible so long as one can quantify the relative probability of any ordering of observed datapoints (or score values).
Formal Sketch of the Key Insight We first sketch analysis from Tibshirani et al. (2019) (which corresponds to viewing CP as an inverted permutation test) followed by the key insight for our current paper’s main result. Their analysis begins by conditioning on the set [^4] For simplicity, we assume all the values are distinct to work with sets rather than multisets; in general, the argument holds with more complex notation using uniform randomness to break ties. of data values (or scores)—that is, on the event $\{Z_{1},...,Z_{n+1}\}=\{z_{1},...,z_{n+1}\}$ denoted by $E_{z}$. To be clear, $E_{z}$ means the unordered “bag” of values is known but not whether $Z_{i}=z_{i}$, and so on; we know that the random variable $Z_{i}$ takes on an observed value in $\{z_{1},...,z_{n+1}\}$, but we do not yet know which value (similarly for the scores $V_{i}$). Given $E_{z}$, the goal is to examine the probability that the test point’s score $V_{n+1}$ was actually $v_{i}$, for each $i\in\{1,...,n+1\}$, which they write as
@@
where $f$ denotes the joint probability density function (PDF). Standard CP proceeds from this statement by substituting $f(z_{\sigma(1)},...,z_{\sigma(n+1)})=f(z_{1},...,z_{n+1})$ via exchangeability, reducing the problem to counting permutations; weighted CP proceeds by factorizing $f$ using weighted exchangeability, with the product of weight functions representing the “weight” given to each permutation $\sigma$; and, CP with pseudo-exchangeable data proceeds similarly to handle a specific instance of potential dependencies (details in Appendix D).
Generally, the key insight in our paper is that simplifying Eq. (4) is only a practical rather than a theoretically necessary step. For intuition on this, note that Eq. (4) makes no assumptions on $f$; it follows from the definition of conditional probability and the law of total probability (LOTP):
$\mathbb{P}\{Z_{n+1}=z_{i}\ |\ E_{z}\}$ $=\frac{p\{Z_{n+1}=z_{i},E_{z}\}}{p\{E_{z}\}}$
@@
That is, the last step uses LOTP in both the numerator and denominator—for instance, the denominator’s sum is over every way the event $E_{z}$ could occur, which amounts to all the possible permutations $\sigma$ of the values. The crux of extending CP to any, possibly non-exchangeable $f$ can thus reduce to computing $f(z_{\sigma(1)},...,z_{\sigma(n+1)})$ directly for all orderings $\sigma$, or factoring $f$ into “dynamic” terms that are tractable to compute (e.g., “weight functions”) and remaining permutation-invariant terms that cancel in the ratio.
@@ 4.2 A General View of Conformal Prediction
Limitations and Benefits of the General View of CP There are two main obstacles to computing Eq. (6) in practice: a computational complexity challenge and an epistemic challenge regarding knowledge about $f$. That is, in the general case for an arbitrary or unfactorized $f$, Eq. (4) requires $\mathcal{O}((n+1)!)$ evaluations of $f$, which quickly becomes intractable. Secondly, we often do not know $f$; indeed, the primary motivation of distribution-free UQ is that very deficiency. Exchangeability, weighted exchangeability (Tibshirani et al., 2019), and related conditions such as pseudo-exchangeability (Fannjiang et al., 2022) introduce distributional assumptions that alleviate these challenges.
@@
One immediate benefit of the general presentation of conformal prediction in Theorem 4.1 is a shift in perspective: because the theorem holds for any joint distribution $f$, for any particular $f$ the question is not whether a CP set with valid coverage exists, it is instead whether we can compute the prediction set in practice. Consequently—from this general perspective—“quasi-exchangeability” definitions are no longer necessary conditions for CP theory; rather, the only role of these definitions is to enable practical tractability by simplifying and eliminating terms from the computation (of Eq. (4))—for example, at one extreme, exchangeability “assumes away” $f$ from the computation entirely. (Appendix D provides further formal details on this point, including how previous CP validity guarantees can be viewed as corollaries to Theorem 4.1.) Specific weighted CP algorithms with corresponding coverage validity guarantees can thus be derived for any (possibly non-exchangeable) distribution. In section 4.3, we outline how to perform such derivations in general and provide the MFCS case as a worked example.
@@
Further Discussion of Tibshirani et al. (2019) This general view of CP is implicit in a close reading of Tibshirani et al. (2019), [^6] We note that the main insight in our paper is also invoked in the following lecture notes (Tibshirani, 2023): https://www.stat.berkeley.edu/~ryantibs/statlearn-s23/lectures/conformal_ds.pdf, which a sharp reader could use to infer Theorem 4.1. However, based on how Tibshirani et al. (2019) has been cited in the CP literature (i.e., primarily as extending CP to standard covariate shift or to weighted exchangeability), we believe this important insight has not been fully appreciated by the community and we hope that our contribution can help correct this oversight. but in that paper the authors did not explicitly state or prove the general result we present in Theorem 4.1. Instead, Tibshirani et al. (2019) premised their main result on their definition of weighted exchangeability, although this condition is not strictly necessary for CP in the broad sense. While we maintain that weighted exchangeability is still a useful definition (see Appendix D.2), subsequent literature has seemingly been led to believe that new definitions of “quasi-exchangeability” are necessary for advancing and delineating the boundaries of CP theory—for example, Fannjiang et al. (2022)’s introduction of pseudo-exchangeability to study the single-step FCS setting. We hope our contribution will lead to wider recognition of the generality of Tibshirani et al. (2019)’s analysis and foster a broadened scope of possibility in the study of CP: in particular, we encourage a shift away from viewing “quasi-exchangeability” conditions as theoretical boundaries to be overcome, and instead towards primarily using such conditions to clearly specify the assumptions granted by a given setting of interest or adopted for practical tractability.
@@
4.3 How to Find Weighted CP Validity Guarantees and Algorithms for MFCS (or Any Data Distribution)